- COM(only control ): Nakano, Guiguo
1/25-2/20: Yend Pcal finalization
- 2/25-3/4: Yend WAB
- 3/6~ : Yend pumpdown 3/6~
- 3/7,8: leak test
2/11-2/14: Xend NAB installation
2/21: Removing FC, Payload release
- 2/23-3/10: Xend Pcal finalization 1
- Periscope installation
- Beam position by IR and Green
- 2/27: ETMX release
- 3/5: ETMX aligned
- 3/5-3/10: Xend Pcal finalization 2
- 3/11,12 Xend WAB finalization
- 3/19,20, Heatlink etc.
- 3/22 Pumping down
- Leak test
- 3/11~ : Night work starts
- 2/18-3/22 DRMI commissioning
- 3/11-4/5 Y arm commissioning
- 3/25-4/5 FPMI commissioning
Report of work
- CRY: FC of ETMX was removed, and cryo payload was released.
- COM: Michelson was locked again!
- Tried to lock PRMI, but PRM was not moving somehow. Actually PRMI was locked for 1 or 2 second.
- SR3 got aligned and green reached to Y end.
- SR2 was also aligned.
- Type B: SRM seems to be healthy.
- DGS: Dolphin connection for PRM and BS were done.
- Type A: needs time for tuning of ITMX.